# Quantifying Multilingual Performance of Large Language Models Across Languages

### v1

@@ 4 Conclusions and Future Work
In this work, we propose a similarity-based evaluation method to measure the LLMs’ performance in various languages quantitatively. The results show that this similarity metric has a clear correlation with the proportion of each language in the pre-training corpus, and can roughly measure the performance ability of the model in each language. In the future, we plan to design more detailed evaluation criteria to measure LLM’s capabilities in each language.
@@
The proposed Language Ranker approach provides an initial quantitative way to analyze LLM performance across languages. However, it has several limitations. First, the similarity metric based on English representations may not fully capture the nuances and complexities of each language’s linguistic properties. Additionally, low-resource languages are likely to exhibit more noise and variance in the similarity scores due to the smaller dataset sizes used for pre-training these languages in LLMs. Furthermore, the method does not account for potential biases or skews that could be present in the multilingual evaluation datasets themselves. The existence of such biases can also introduce noise in the resulting rankings of language abilities for different LLMs.
@@ Appendix A Appendix
A.1 Related Work
Multilingual Language Model.  The imbalance distribution of training corpus in different languages leads to the bias of LLM towards some common languages such as English Blasi et al. (2021). Some approaches employ multilingual language modeling to alleviate the phenomenon Shen et al. (2024); Kalyan et al. (2021); Conneau et al. (2019). These studies show the importance of strengthening the cross-lingual capabilities of the pre-trained model. Schäfer et al. (2024) found that the presence of a primary language in the training process of LLMs can improve the performance of low-resource languages and lead to a more consistent representation of LLMs in different languages. Liu et al. (2024) found that for English-centric LLMs, although translation into English helps improve the performance of NLP tasks, it is not the best choice for all situations.
A.2 Ranking Result For LLMs
We give the similarity scores of the four LLMs used in the experiment on 18 common languages. Results are shown in following tables.
@@ A.2 Ranking Result For LLMs
German 0.581 Polish 0.534
French 0.592 Portuguese 0.598
Swedish 0.531 Vietnamese 0.529
Chinese 0.446 Ukrainian 0.551
Spanish 0.616 Korean 0.199
Russian 0.589 Catalan 0.582
Dutch 0.569 Serbian 0.555
Italian 0.567 Indonesian 0.577
Japanese 0.194 Czech 0.587
@@
German 0.571 Polish 0.487
French 0.546 Portuguese 0.535
Swedish 0.494 Vietnamese 0.456
Chinese 0.471 Ukrainian 0.484
Spanish 0.537 Korean 0.338
Russian 0.531 Catalan 0.492
Dutch 0.516 Serbian 0.472
Italian 0.522 Indonesian 0.499
Japanese 0.328 Czech 0.512
@@
German 0.516 Polish 0.421
French 0.503 Portuguese 0.482
Swedish 0.435 Vietnamese 0.392
Chinese 0.399 Ukrainian 0.437
Spanish 0.499 Korean 0.261
Russian 0.494 Catalan 0.445
Dutch 0.460 Serbian 0.408
Italian 0.466 Indonesian 0.421
Japanese 0.248 Czech 0.459
@@
German 0.642 Polish 0.596
French 0.634 Portuguese 0.625
Swedish 0.603 Vietnamese 0.584
Chinese 0.608 Ukrainian 0.597
Spanish 0.638 Korean 0.481
Russian 0.634 Catalan 0.601
Dutch 0.612 Serbian 0.588
Italian 0.615 Indonesian 0.597
Japanese 0.457 Czech 0.611
@@
A.3 Performance Comparison of Qwen1.5 of Different Sizes
The following two figures show the performance of Qwen1.5 in different sizes for different languages. Figure 4 shows the result of five high-resource languages. Figure 5 shows the result of five low-resource languages. From Figure 4, we observe that the performance of 0.5B is the best, while 7B performs the worst. Figure 5 shows the opposite result. It can also be found that the performance variance in low-resource languages is much greater than the performance in high-resource languages.
Figure 4: Result for five high-resource languages.
Figure 5: Result for five low-resource languages.

### v2

@@ 6 Conclusions and Future Work
In this work, we propose the Language Ranker to evaluate the performance of LLMs across diverse languages by comparing their internal representations to English. The results show that high-resource languages show higher similarity scores with English, while low-resource languages have lower scores, validating the effectiveness of our metric in assessing language performance. Besides, there is a strong correlation between the performance of LLMs in different languages and the proportion of those languages in the pre-training corpus. Further, results indicate that high-resource languages are more evenly distributed in the embedding space, whereas low-resource languages tend to be narrowly clustered. In the future, we plan to design more comprehensive benchmarks to measure LLM’s capabilities in different languages.
@@
The proposed Language Ranker approach provides an initial quantitative way to analyze LLM performance across languages. We acknowledge that the language ranker method we proposed offers only a rough measurement. While our findings indicate a correlation between the similarity scores and the proportion of each language in the pre-training dataset, these scores alone are not sufficient to precisely measure the exact proportions. Our intent was to provide an initial quantitative approach to explore this relationship, and we recognize the need for more comprehensive methods and additional metrics to accurately assess the impact of pre-training data distribution across languages. Furthermore, the method does not account for potential biases or skews that could be present in the multilingual evaluation datasets themselves. The existence of such biases can also introduce noise in the resulting rankings of language abilities for different LLMs.
@@ Appendix A Appendix
A.1 Ranking Result For LLMs
We give the similarity scores of the four LLMs used in the experiment on 18 high-resource languages. Results are shown in the following tables.
@@ A.1 Ranking Result For LLMs
German 0.723 Western Frisian 0.378
French 0.737 Tamil 0.347
Spanish 0.768 Gujarati 0.313
Italian 0.706 Kurdish 0.308
Russian 0.734 Pashto 0.284
Dutch 0.709 Assamese 0.260
Polish 0.664 Central Khmer 0.240
Malay 0.651 Panjabi 0.218
Swedish 0.661 Amharic 0.202
@@
German 0.719 Western Frisian 0.443
French 0.691 Tamil 0.420
Spanish 0.683 Gujarati 0.433
Italian 0.662 Kurdish 0.358
Russian 0.674 Pashto 0.362
Dutch 0.658 Assamese 0.396
Polish 0.618 Central Khmer 0.330
Malay 0.615 Panjabi 0.379
Swedish 0.629 Amharic 0.298
@@
German 0.639 Western Frisian 0.346
French 0.623 Tamil 0.279
Spanish 0.616 Gujarati 0.270
Italian 0.571 Kurdish 0.262
Russian 0.611 Pashto 0.267
Dutch 0.566 Assamese 0.276
Polish 0.514 Central Khmer 0.252
Malay 0.497 Panjabi 0.213
Swedish 0.532 Amharic 0.191
@@
German 0.805 Western Frisian 0.441
French 0.793 Tamil 0.510
Spanish 0.800 Gujarati 0.469
Italian 0.773 Kurdish 0.436
Russian 0.794 Pashto 0.448
Dutch 0.773 Assamese 0.507
Polish 0.752 Central Khmer 0.407
Malay 0.730 Panjabi 0.385
Swedish 0.759 Amharic 0.470
@@
Figure 6: Simlarity scores curves of LlaMa-13b and Qwen-13b.
Figure 7: Simlarity scores curves of Bloom-3b and Bloom-7b.
A.2 Details of experiment in Table 4 and Table 3
We selected data from the Tatoeba-Challenge repository [^2] https://github.com/Helsinki-NLP/Tatoeba-Challenge/tree/master/data. Since the number of samples for some low-resource language pairs is small, we extracted 100 samples for each language pair. If there are less than 100, we extracted all samples and extracted them according to the test-dev-train priority.
High-High Similarity Score High-Low Similarity Score Low-Low Similarity Score
English-German 0.64 German-Silesian 0.43 Azerbaijani-Turkmen 0.51
Italian-French 0.60 French-Erzya 0.30 Hungarian-Yiddish 0.18
German-French 0.62 Italian-Romany 0.25 Kab-SMT 0.39
French-Chinese 0.57 Italian-Uighur 0.24 Mari-Tatar 0.46
Table 10: Similarity score of different language pairs of mistral-7b.
High-High Similarity Score High-Low Similarity Score Low-Low Similarity Score
English-German 0.81 German-Silesian 0.58 Azerbaijani-Turkmen 0.67
Italian-French 0.75 French-Erzya 0.51 Hungarian-Yiddish 0.39
German-French 0.80 Italian-Romany 0.53 Kab-SMT 0.55
French-Chinese 0.71 Italian-Uighur 0.49 Mari-Tatar 0.55
Table 11: Similarity score of different language pairs of qwen-7b.
A.3 Performance of LLMs with Other Sizes
A.3.1 LlaMa-13b and Qwen-13b
We also explore the performance of the similarity metric in LLM of 13B parameters, We use LlaMa-13b and Qwen-13b to display the results. From Figure 6 we can observe that the partial order results of LLM-13B are roughly the same as those of LLM-7B, and there is a clear gap between high-resource languages and low-resource languages.
A.3.2 Bloom-3b and Bloom-7b
Figure 7 shows the result of Bloom-3b and Bloom-7b. Except for the last few layers of the model, there are obvious differences between high-resource languages and low-resource languages which are similar to the above LLMs, while there are smaller differences within the same category of languages. The score is higher than LlaMa, Gemma, Mistral, and Qwen.